OpenStudy (anonymous):

If m = 65, m = 2x, and m = 3x – 10, what is the value of x? 15 65 13 35

jhonyy9 (jhonyy9):

this mean that 65=2x=3x-10 can you continue ?

jhonyy9 (jhonyy9):

65=x-10

jhonyy9 (jhonyy9):

than x=?

OpenStudy (anonymous):

75

jhonyy9 (jhonyy9):

so i think you have missed anything

jhonyy9 (jhonyy9):

ye

jhonyy9 (jhonyy9):

yes 75

OpenStudy (anonymous):

but 75 is not an option for my problem :(

jhonyy9 (jhonyy9):

yes i see it the same

OpenStudy (anonymous):

So what do I have to do now ???

jhonyy9 (jhonyy9):

65=2x 65=3x-10 --- 3x=75 2x=65 3x=75 2x=65 ------- subtract both sides 3x-2x=75-65 x=10

jhonyy9 (jhonyy9):

or assuming sorry i think there is missed anything

jhonyy9 (jhonyy9):

or these choises not are completly

jhonyy9 (jhonyy9):

so try it getting 65 using 65=2x and 65=3x-10 using for x these choises

jhonyy9 (jhonyy9):

i think not is possibill

jhonyy9 (jhonyy9):

65=3x-10 75=3x x=25